OpenStudy (anonymous):

use subtraction 4x−3y=11 10x−3y=41

OpenStudy (anonymous):

4x−3y=11 10x−3y=41 10x−3y -(4x−3y)=41 - 11 10x - 3y -4x + 3y = 30 6x = 30 find x and y

OpenStudy (anonymous):

there are question marks everywhere??

OpenStudy (anonymous):

reload the page. and if still happens please tell